# On the Viability of Open-Source Financial Rails: Economic Security of Permissionless Consensus

### v1

@@ Abstract
Bitcoin demonstrated the possibility of a financial ledger that operates without the need for a trusted central authority. However, concerns persist regarding its security and considerable energy consumption. We assess the consensus protocols that underpin Bitcoin’s functionality, questioning whether they can ensure economically meaningful security while maintaining a permissionless design that allows free entry of operators. We answer this affirmatively by constructing a protocol that guarantees economic security and preserves Bitcoin’s permissionless design. This protocol’s security does not depend on monetary payments to miners or immense electricity consumption, which our analysis suggests are ineffective. Our framework integrates economic theory with distributed systems theory, and highlights the role of the protocol’s user community.
@@ 2.4 Definition of economic security
Conversly, if an attacker deceives a merchant who follows the protocol, then consistency is violated.
We combine the distributed systems model with an economic model to define economic security. Classic results in the distributed systems literature prove consistency and liveness of protocols given a bound on the number of faulty nodes. For example, classic consensus protocols such as Paxos (Lamport,, 2001) and PBFT (Castro et al.,, 1999) guarantee consistency and liveness when more than 2/3 of the nodes are honest. [^28] These classic consensus protocols are permissioned, that is, designed for settings where a central designer authorizes each node (for example, a firm operating multiple data centers). In such settings, faults may occur due to hardware outages or software errors, and it is natural to assume that faulty behavior of many nodes is rare. However, consistency may be violated if a majority of nodes are faulty.
@@
3 Attacks on Bitcoin’s Nakamoto Protocol: Shortcomings of Incentives
We evaluate Nakamoto’s economic security in two economic environments, which are chosen to highlight the deficiencies of incentives in securing permissionless protocols. First, the protocol lacks a mechanism to penalize attackers. Section 3.2 describes how the protocol can be attacked, and Section 3.3 details the design choices within the protocol that prevent it from identifying and punishing attacks. Second, even if an attack prompts a response from outside the protocol (e.g., an exchange rate collapse), this response negatively impacts all miners regardless of their involvement in the attack. Section 3.4 shows that if miners are small, the tragedy of the commons can arise.
The results suggest that Bitcoin’s security does not stem from miners’ monetary incentives. We provide an alternative explanation for the security of Bitcoin in Section 3.5. Additional explanations are discussed in Section D.
@@ 3.1 Background: Bitcoin’s Nakamoto protocol
The ledger is saved as a blockchain that grows by appending new blocks over time. A blockchain is an ordered sequence of blocks $({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n})$, where ${\sf B}_{0}$ is the protocol’s genesis block. Each block ${\sf B}_{\ell}$ contains transaction data and a reference to the preceding block ${\sf B}_{\ell-1}$. Thus, the final block ${\sf B}_{n}$ uniquely identifies the entire chain, namely, $\sf Chain({\sf B}_{n})=({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n})$. We say that block ${\sf B}_{n-k}$ is $k$-deep in $\sf Chain({\sf B}_{n})$. Let $h({\sf B}_{\ell})=\ell$ denote the height of the block, defined as the number of blocks in $\sf Chain({\sf B}_{\ell})$ minus one.
The view of node $i\in N$ is the information available to node $i$ and includes $i$’s local history and any block that $i$ received. A node may receive multiple blocks of the same height that cannot be in the same chain.
@@
The Nakamoto protocol itself does not specify when transactions are finalized. In common usage, finality is parameterized by the required confirmation depth $k$. A transaction is considered finalized by node $i$ if the transaction is within a block at least $k$-deep in the longest chain in $i$’s view. That is, if $\sf Chain({\sf B}_{n})=({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n})$ is (one of) the longest chains in the view of node $i$, then $i$’s local copy of the ledger is ${\sf Log}_{i}=({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n-k+1})$.
@@ 3.2 Economic security of Bitcoin under the rental model
We consider a stylized economic environment where miners can freely enter and exit. Several papers have explored a similar model and argued that an attacker can violate the consistency of Nakamoto at a low cost (Budish, 2022, Tabarrok, 2019, Auer, 2019, Moroz et al., 2020, Gans and Halaburda, 2024). We provide a similar result in our framework, showing that protocol users cannot defend against these attacks and clarifying the protocol’s design choices that enable such attacks.
Miners choose their computational power and may flexibly adjust it (for example, by renting capacity from a cloud computing service). Computing costs are linear: each miner pays $c$ USD per attempt to compute a hash. Each hash provides the miner a $1/D$ chance of mining a block. Each block includes a reward of $p_{\sf B}$. Miners are purely profit-motivated. The attacker has the same capabilities and costs as other miners.
@@
The following theorem states that Nakamoto’s economic security in this economic environment is zero, as the attacker can violate consistency at zero cost. Moreover, this holds even if the merchant uses additional security measures, such as requiring a longer confirmation period or monitoring network activity.
@@
The attacker deceives the merchant by creating a conflicting ledger: a blockchain in which the transaction crediting the merchant does not appear. Because the protocol is permissionless, such a blockchain can be created by operating a set of miners and giving them malicious inputs. [^34] The miners employed by the attacker can even run the same code as honest miners. The attacker can direct its miners to run the protocol honestly in a separate environment where the miners never receive the merchant’s transaction or any blocks in a chain that includes it. The attacker waits for the merchant to finalize the transaction before launching the attack, and therefore, the merchant cannot protect itself by monitoring. The attacker exploits the longest-chain rule to get honest nodes to adopt the conflicting ledger. If the attacker’s blockchain becomes longer than the honest chain, all honest nodes adopt its conflicting chain, violating consistency.
The proof uses a similar approach as Budish, (2022), Auer, (2019), and Gans and Halaburda, (2024), and is included in Appendix B for completeness.353535 Budish, (2022) and Auer, (2019) show that Nakamoto has positive economic security if the attacker cannot recoup the mining rewards or if the attacker has higher mining costs. Gans and Halaburda, (2024) show that the net cost of the attack can be negative because the attacker can affect the block reward. We discuss such factors in Section D. The key argument in the proof is that the protocol fails to recognize that an attack has occurred and treats blocks mined by the attacker as if they were mined by honest miners. In particular, once the attacker’s chain is adopted, the attacker becomes the “honest miner” who created the agreed ledger and is paid the mining rewards. Since honest miners do not lose from mining, neither does the attacker.
@@
3.3 Nakamoto’s design choices that enable attacks
To understand why the protocol permits such attacks, it is useful to examine its design and the rationale behind its longest-chain rule.
@@ 3.4 Bribery model and the tragedy of the commons
The bribery model has a continuum of miners, each controlling an infinitesimal unit of computational power.424242 Contemporaneously and independently of this work, Newman, (2023) analyzes a model with finitely many miners and shows that small miners are unlikely to be pivotal to an attack’s success. Let $\psi_{A}$ denote the normalized harm per infinitesimal miner. Miners will participate in the attack if it is in their selfish best interest, given the payments they stand to receive from the protocol, the bribes from the miner, and the harm they may suffer from an attack.
@@ 3.5 What deters attacks in practice?
Despite the possibility of attacks, major cryptocurrencies like Bitcoin and Ethereum [^45] Ethereum operated a variant of the Nakamoto protocol until 2022. have remained secure. At the same time, numerous attacks on multiple cryptocurrencies that operate the Nakamoto protocol show that the Nakamoto protocol is indeed vulnerable to attacks in practice. [^46] See Shanaev et al., (2019) for a list of attacks.
We discuss several explanations for the absence of attacks on Bitcoin in Appendix D and highlight one explanation in particular: the community’s ability to correct the ledger after an attack. Recall that the protocol provides a record-keeping service: maintaining an agreed-upon ledger. If the community agrees that the ledger was corrupted, the community may override the protocol to change it. In blockchain terminology, the community can “fork” the protocol by directing all miners to ignore the attack chain and extend the honest chain. Doing so reverts any effect of the attack, leaving the ledger unchanged by the attack. [^47] Such corrections also occur in traditional record-keeping services. For example, the New York Stock Exchange canceled trades that resulted from a software glitch (https://link.cnbc.com/public/35649502, accessed June 2024). Moreover, the community response imposes a significant risk to the attacker of having to pay all the attack’s costs only to see it undone.
Such community forks are not without precedent. A notable example is the Ethereum community’s decision to fork the Ethereum ledger in order to return funds stolen from the DAO attack.484848 https://blog.ethereum.org/2016/07/20/hard-fork-completed, accessed February 2024. Such measures, however, require the community to be willing to override the protocol and able to agree on the correct ledger. While the Ethereum blockchain community was able to override the protocol, the Ethereum Classic blockchain rejects any community interventions. [^49] https://ethereumclassic.org/why-classic/code-is-law, accessed February 2024. Anecdotally, Ethereum did not suffer any consistency attacks, whereas Ethereum Classic suffered multiple such attacks. [^50] https://www.coinbase.com/blog/deep-chain-reorganization-detected-on-ethereum-classic-etc, accessed February 2024.
However, the community response is not a silver bullet. To undo the attack, the community must reach agreement on the correct ledger. In other words, the community needs to solve the same consensus problem that the protocol attempts to solve. In addition, the attacker may employ a different attack to deceive both the protocol and the community.
@@
4 An Economically Secure Permissionless Consensus Protocol
@@
We show that a permissionless protocol can be economically secure by constructing such a protocol. This protocol is a variation of the Nakamoto protocol, and we name it the Stubborn Nakamoto protocol. We prove that it is infinitely economically secure (i.e., always consistent) in the same communication model as Nakamoto (namely, synchrony with late joiners). The Stubborn Nakamoto protocol provides economic security while maintaining the permissionless properties of the Nakamoto protocol; that is, miners in the Stubborn Nakamoto protocol are anonymous and can enter and exit without restriction.
The security of the Stubborn Nakamoto protocol does not rely on monetary incentives or mining rewards. Instead, the protocol’s security comes from communication between the nodes. Loosely speaking, similarly to how a merchant may call the bank to verify a payment, nodes in the Stubborn Nakamoto protocol communicate with the rest of the network. An attacker cannot deceive protocol participants because a node that receives conflicting records can communicate with the rest of the network and detect the conflict.
The protocol is always consistent and requires essentially the same honest majority assumption as Nakamoto for liveness. No protocol can satisfy consistency and liveness without honest majority (Pass and Shi, 2017a, ; Lewis-Pye and Roughgarden,, 2023). Since Stubborn Nakamoto is always consistent, it may lose liveness under some corrupt majority attacks. Namely, the protocol directs nodes to halt (stop mining and finalizing blocks) if an attack could otherwise compromise consistency. We show that the protocol can recover from such attacks with some external help and that some external help is necessary for recovery.
@@
Section 4.1 defines the Stubborn Nakamoto protocol, and we prove its security and liveness in Section 4.2. Section 4.3 details how the Stubborn Nakamoto protocol can recover from an attack. Section 4.4 examines the necessary role of the community.
@@ 4.1 The Stubborn Nakamoto protocol
The Stubborn Nakamoto protocol requires the same communication assumptions as Bitcoin’s Nakamoto: synchronous communication with late joiners. It also shares the same setup assumption as Bitcoin’s Nakamoto: all nodes know the genesis block and the protocol. Stubborn Nakamoto is parameterized by a confirmation depth $k$.
We say that node $i$ confirms block ${\sf B}$ at time $t$ if, according to the local view of node $i$, at time $t$ block ${\sf B}$ becomes at least $k$ blocks deep in some longest chain. That is, $i$ confirms block ${\sf B}$ when it would have become newly finalized under Nakamoto. We say that block ${\sf B}$ conflicts with block ${\sf B}^{\prime}\neq{\sf B}$ if the two blocks are of the same height, and thus there is no chain that contains them both. We say that block ${\sf B}$ is first-confirmed in node $i$’s view if at some previous time $t$ node $i$ confirmed block ${\sf B}$ and at time $t$ node $i$ did not confirm any block ${\sf B}^{\prime}\neq{\sf B}$ that conflicts with ${\sf B}$.

### v2

@@ Abstract
Bitcoin demonstrated the possibility of a financial ledger that operates without the need for a trusted central authority. However, concerns persist regarding its security and considerable energy consumption. We assess the consensus protocols that underpin Bitcoin’s functionality, questioning whether they can ensure economically meaningful security while maintaining a permissionless design that allows free entry of operators. We answer this affirmatively by constructing a protocol that guarantees economic security and preserves Bitcoin’s permissionless design. This protocol’s security does not depend on monetary payments to miners or immense electricity consumption, which our analysis suggests are ineffective. Our framework integrates economic theory with distributed systems theory, and formalizes the role of the protocol’s user community.
@@ 1 Introduction
We approach this problem as an economic design problem, asking whether it is possible to design a protocol that has the free entry properties that are provided by a permissionless design while also providing a security guarantee that is economically meaningful.
@@ 2.4 Definition of economic security
Conversely, if an attacker deceives a merchant who follows the protocol, then consistency is violated.
We combine the distributed systems model with an economic model to define economic security. Classic results in the distributed systems literature prove the consistency and liveness of protocols given a bound on the number of faulty nodes. For example, classic consensus protocols such as Paxos (Lamport,, 2001) and PBFT (Castro et al.,, 1999) guarantee consistency and liveness when more than 2/3 of the nodes are honest. [^28] These classic consensus protocols are permissioned, that is, designed for settings where a central designer authorizes each node (for example, a firm operating multiple data centers). In such settings, faults may occur due to hardware outages or software errors, and it is natural to assume that faulty behavior of many nodes is rare. However, consistency may be violated if a majority of nodes are faulty.
@@
3 Incentives in Permissionless Protocols: Analysis of Bitcoin’s Nakamoto Protocol
We formally demonstrate challenges in securing permissionless protocols by applying our definition of economic security to Bitcoin’s Nakamoto protocol. Our analysis focuses on two distinct economic environments, chosen to highlight the inherent incentive deficiencies. First, the protocol lacks a mechanism to penalize attackers. Section 3.2 describes how the protocol can be attacked, and Section 3.3 details the design limitations that prevent it from detecting and punishing attacks. Second, even if an attack prompts a response from outside the protocol (e.g., an exchange rate collapse), this response negatively impacts all miners regardless of their involvement in the attack. Section 3.4 shows that if miners are small, the tragedy of the commons can arise.
The results suggest that Bitcoin’s security does not stem from miners’ monetary incentives. We provide an alternative explanation for the security of Bitcoin in Section 3.5. Other considerations are discussed in Appendix D.
@@ 3.1 Background: Bitcoin’s Nakamoto protocol
The ledger is held in the format of a blockchain, which grows over time as new blocks are appended. A blockchain is an ordered sequence of blocks $({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n})$, where ${\sf B}_{0}$ is the protocol’s genesis block. Each block ${\sf B}_{\ell}$ contains transaction data and a reference to the preceding block ${\sf B}_{\ell-1}$. A block ${\sf B}_{n}$ uniquely identifies a chain of preceding blocks, and we write $\sf Chain({\sf B}_{n})=({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n})$. We say that block ${\sf B}_{n-k+1}$ is $k$-deep in $\sf Chain({\sf B}_{n})$. Let $h({\sf B}_{\ell})=\ell$ denote the height of the block, defined as the number of blocks in $\sf Chain({\sf B}_{\ell})$ minus one.
The view of a node $i\in N$ encompasses all the information available to node $i$, including its local history and any blocks it has received. A node may receive multiple blocks of the same height, which cannot belong to the same chain.
@@
The Nakamoto protocol itself does not specify when transactions are finalized. In common usage, [^33] The parameter $k$ is commonly called the “confirmation depth”. A common value for $k$ is $6$ blocks. For example, see https://coinmarketcap.com/academy/article/how-long-does-a-bitcoin-transaction-take. finality is parameterized by a required block depth $k$, where a transaction is considered finalized by node $i$ if the transaction is within a block at least $k$-deep in the longest chain in $i$’s view. That is, if $\sf Chain({\sf B}_{n})=({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n})$ is (one of) the longest chains in the view of node $i$, then $i$’s local copy of the ledger is ${\sf Log}_{i}=({\sf B}_{0},{\sf B}_{1},\dots,{\sf B}_{n-k+1})$.
@@ 3.2 Economic security of Bitcoin under the rental model
We consider a stylized economic environment where miners can freely enter and exit. Several papers have explored a similar model and argued that an attacker can violate the consistency of Nakamoto at a low cost (Budish, 2022, Tabarrok, 2019, Auer, 2019, Moroz et al., 2020, Gans and Halaburda, 2024). We provide a similar result in our framework, showing that protocol users cannot defend against these attacks and clarifying why the protocol’s design enables such attacks.
In the rental model, miners choose their computational power and may flexibly adjust it (for example, by renting capacity from a cloud computing service). Computing costs are linear: each miner pays $c$ per attempt to compute a hash. Each hash provides the miner a $1/D$ chance of mining a block. Each block includes a reward of $p_{\sf B}$. Miners are purely profit-motivated. The attacker has the same capabilities and costs as other miners.
@@
The following theorem states that Nakamoto’s economic security in this economic environment is zero, as the attacker can violate consistency at zero cost. Moreover, this holds even if the merchant uses additional security measures, such as requiring additional verifications before finalizing transactions or monitoring network activity.
@@
The attacker deceives the merchant by creating a conflicting ledger: a blockchain in which the transaction crediting the merchant is absent. This is possible because of the permissionless nature of the protocol, as the attacker can construct such a blockchain by operating a set of miners and providing them with malicious inputs. [^35] The attacker’s miners can even run the same code as honest miners. However, the attacker directs its miners to operate in a separate environment where they never receive the merchant’s transaction or any blocks from a chain that includes it. The attacker waits for the merchant to finalize the transaction before initiating the attack, rendering any monitoring by the merchant ineffective. The critical phase of the attack involves persuading honest nodes to adopt the conflicting ledger. This is achievable because of the longest chain rule: if the attacker’s blockchain becomes longer than the honest chain, all honest nodes will adopt the conflicting chain, thereby violating consistency.
The proof uses a similar approach as Budish, (2022), Auer, (2019), and Gans and Halaburda, (2024), and is included in Appendix B for completeness.363636 Budish, (2022) and Auer, (2019) show that Nakamoto has positive economic security if the attacker cannot recoup the mining rewards or if the attacker has higher mining costs. Gans and Halaburda, (2024) show that the net cost of the attack can be negative because the attacker can affect the block reward. We discuss such factors in Section D. The key argument in the proof is that the protocol fails to recognize that an attack has occurred and treats blocks mined by the attacker as if they were mined by honest miners. In particular, once the attacker’s chain is adopted, the attacker becomes the “honest miner” who created the agreed ledger and is paid the mining rewards. Since honest miners do not lose from mining, neither does the attacker.
@@
3.3 Why Nakamoto fails to detect attacks
To understand why the Nakamoto protocol permits such attacks, it is useful to examine its design and the rationale behind its longest-chain rule.
@@ 3.4 Bribery model and the tragedy of the commons
The bribery model has a continuum of miners, each controlling an infinitesimal unit of computational power.434343 Contemporaneously and independently of this work, Newman, (2023) analyzes a model with finitely many miners and shows that small miners are unlikely to be pivotal to an attack’s success. Let $\psi_{A}$ denote the normalized harm per infinitesimal miner. Miners will participate in the attack if it is in their selfish best interest, given the payments they stand to receive from the protocol, the bribes from the miner, and the harm they may suffer from an attack.
@@ 3.5 What deters attacks in practice?
Despite the possibility of attacks, major cryptocurrencies like Bitcoin and Ethereum [^46] Ethereum used the Nakamoto protocol until 2022. have remained secure. At the same time, numerous attacks on multiple cryptocurrencies that operate the Nakamoto protocol show that the Nakamoto protocol is indeed vulnerable to attacks in practice (Shanaev et al.,, 2019).
We discuss several explanations for the absence of attacks on Bitcoin in Appendix D and highlight one explanation in particular: the community’s ability to correct the ledger after an attack, which we refer to as a community response. Recall that the protocol provides a record-keeping service: maintaining an agreed-upon ledger. If the community agrees that the ledger was corrupted, the community may override the protocol to change it. In blockchain terminology, the community can “fork” the protocol by directing all miners to ignore the attack chain and extend the honest chain. Doing so reverts any effect of the attack, leaving the ledger unchanged by the attack. [^47] Such corrections also occur in traditional record-keeping services. For example, the New York Stock Exchange canceled trades that resulted from a software glitch (CNBC,, 2024). Moreover, the community response imposes a significant risk to the attacker of having to pay all the attack’s costs only to see it undone.
[…]
However, the community response is not a silver bullet. To undo the attack, the community must reach an agreement on the correct ledger. In other words, the community needs to solve the same consensus problem that the protocol attempts to solve. Moreover, the attacker may employ a different attack to deceive both the protocol and the community. Without a formal model of the community’s decision-making process, it is difficult to assess the community response.
@@
4 Possibility of a Permissionless and Economically Secure Consensus Protocol
@@
We now consider the broader question of protocol design: Can a permissionless [^49] A common alternative approach is to modify the protocol’s permissionless properties. For example, Proof of Stake (PoS) protocols require nodes to register an identifier and associate it with an account. We discuss such protocols in Section 5. protocol achieve economic security? This section answers affirmatively. We show that a permissionless protocol can be economically secure by constructing such a protocol. This protocol is a variation of the Nakamoto protocol, and we name it Stubborn Nakamoto protocol. We prove that it is infinitely economically secure – that is, even an attacker with unlimited resources cannot violate consistency – under the same communication assumptions required by Bitcoin’s Nakamoto, namely, synchrony with late joiners. The Stubborn Nakamoto protocol preserves the permissionless properties of the Nakamoto protocol while ensuring economic security. Specifically, miners remain anonymous and can enter or exit the system without restriction.
The design approach underlying the Stubborn Nakamoto protocol is motivated by the analysis in Section 3. The negative results in Section 3 demonstrate that the inability to detect attacks compromises the economic security of Bitcoin’s Nakamoto protocol, whereas the effectiveness of the community response highlights the power of attack detection. Accordingly, the Stubborn Nakamoto protocol is designed to enable nodes to detect potential attacks.
The protocol can be interpreted as incorporating the community response which restores the correct ledger following an attack. A straightforward design might require nodes to communicate and reestablish the correct ledger after an attack. The Stubborn Nakamoto protocol takes a more efficient approach: if a node is assured that the community response will restore the correct ledger, it can immediately adjust its local copy to reflect that correct ledger. (For example, a node can reject an attack chain if it is confident that all honest nodes will likewise reject it.) The Stubborn Nakamoto protocol provides nodes with sufficient knowledge about the behavior of other nodes, ensuring that they maintain a consistent ledger.
This analysis in this section examines the role of the “community” – that is, external inputs beyond the protocol. Some external input is unavoidable. For instance, the initialization of any protocol requires an external source (e.g., setting the genesis block in Bitcoin’s Nakamoto protocol), and late joiners must rely on an external input to connect to the protocol (e.g., distinguishing Bitcon’s chain from Dogecoin’s chain). However, reliance on such external inputs should be minimized.505050 Reliance on the community to resolve challenges is often challenging to formally model and analyze. Moreover, reliance on out-of-protocol inputs can introduce fundamental issues. First, the primary objective of the protocol is to establish consensus, so it should not depend on the community’s ability to achieve consensus (conversely, if the protocol cannot establish consensus, why should the community be able to do so?). Second, reliance on external inputs may undermine the protocol’s permissionless properties, potentially giving control over the system to an out-of-protocol entity. This section rigorously evaluates the necessary role of the community, formalizing it as a recovery oracle. We prove that some external dependence is necessary for recovery after attacks, and design the Stubborn Nakamoto protocol to minimize this reliance in two key ways. First, the requirements from the recovery oracle can be practically implemented. Second, the Recovery Oracle should not hold excessive control over the system. The Stubborn Nakamoto protocol is thus designed to minimize both: the recovery oracle is simple to implement, invoked only for attack recovery, and holds minimal influence over the system.
@@
Section 4.1 defines the Stubborn Nakamoto protocol, and we prove its security and liveness for online nodes in Section 4.2. Section 4.3 defines the protocol for late joiners and proves the corresponding guarantees. Section 4.4 details how the Stubborn Nakamoto protocol can recover from an attack and formalizes the role of the community in the recovery. Section 4.5 examines the necessary role of the community.
@@ 4.1 The Stubborn Nakamoto protocol
In contrast to the analysis of Bitcoin’s Nakamoto protocol in Section 3, the definition and analysis of Stubborn Nakamoto protocol require careful consideration of each node’s knowledge, and in particular each node’s knowledge of the state of other nodes. Before giving the definition of the Stubborn Nakamoto protocol, we introduce notation and definitions that represent the knowledge held by a node. Note that this section defines the Stubborn Nakamoto protocol for online nodes. The protocol for late joiners is defined in Section 4.3.
The Stubborn Nakamoto protocol makes the same communication assumptions as Bitcoin’s Nakamoto protocol, namely synchronous communication with late joiners. We also assume “implicit echoing”, namely, if an honest node [^51] We assume that the implicit echoing assumption holds for all nodes, including observer nodes that do not mine. Sridhar et al., (2024) show the necessity of this assumption. receives a message at time $t$, then every honest node receives the message by time $t+\Delta$. Stubborn Nakamoto also shares the same setup assumptions as Bitcoin’s Nakamoto protocol: all nodes are assumed to have knowledge of the genesis block and the protocol’s rules. The protocol is parameterized by a confirmation depth $k$.
Stubborn Nakamoto uses the blockchain data format. The local view of node $i$ includes all the blocks received by $i$, as well as additional information recorded in the form of block labels. We write ${\sf B}\in{\sf View}_{i}^{t}$ to denote that node $i$ received block ${\sf B}$ by time $t$. Each block label is encoded as a subset of blocks in view. We describe below how a node assigns each label given its local view. Labels are never removed. [^52] For example, we have $\mathcal{C}_{i}^{t}\subseteq\mathcal{C}_{i}^{\tau}$ for $\tau\geq t$ (and analogously for all other block labels defined below).
A block ${\sf B}$ conflicts with block ${\sf B}^{\prime}$, denoted ${\sf B}\not\sim{\sf B}^{\prime}$, if there is no chain that contains both ${\sf B},{\sf B}^{\prime}$. That is, neither $\sf Chain(B)\subseteq\sf Chain({\sf B}^{\prime})$ nor $\sf Chain(B^{\prime})\subseteq\sf Chain({\sf B})$. Given a block ${\sf B}$ and a set of blocks $\mathcal{A}$, if $\nexists{\sf B}^{\prime}\in\mathcal{A}$ such that ${\sf B}\not\sim{\sf B}^{\prime}$ we say that ${\sf B}$ does not conflict with $\mathcal{A}$, written ${\sf B}\sim\mathcal{A}$. We write ${\sf B}\not\sim\mathcal{A}$ if $\exists{\sf B}^{\prime}\in\mathcal{A}$ such that ${\sf B}\not\sim{\sf B}^{\prime}$.
The protocol may instruct a node to halt if the node detects a potential conflict between its local view and the local view of another honest node. When a node halts it stops its attempts to mine any new blocks and stops certifying any new blocks (as described below). If the node has already certified some blocks, it completes the steps below to determine whether these blocks should be finalized.
The Stubborn Nakamoto protocol relies on the following definition, which specifies the first step a node takes toward finalizing a block. The conditions for this first step closely resemble the conditions for finalizing a block in the Nakamoto protocol. However, unlike the Nakamoto protocol, the required confirmation depth $k$ is explicitly defined within the protocol, and all nodes use the same confirmation depth $k$.
@@
Node $i$ certifies block ${\sf B}$ at time $t$, written ${\sf B}\in\mathcal{C}_{i}^{t}$, if at time $t$ there is a chain $({\sf B}_{0},\dots,{\sf B},\dots,\bar{{\sf B}})$ in ${\sf View}_{t}^{i}$ such that
  1. 1.
    Block ${\sf B}$ is at least $k$-deep in the chain $\sf Chain(\bar{{\sf B}})$, i.e., $h(\bar{{\sf B}})-h({\sf B})+1\geq k$.
  2. 2.
    The chain $\sf Chain(\bar{{\sf B}})$ is a longest chain in view, meaning $h(\bar{{\sf B}})\geq h(\tilde{{\sf B}})$ for any $\tilde{{\sf B}}\in{{\sf View}}_{i}^{t}$.
Additionally, if in the view of node $i$ at time $t$ there is no certified block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$, then ${\sf B}$ is also clear-certified, denoted ${\sf B}\in\mathcal{CC}_{i}^{t}$. [^53] If two conflicting blocks both become certified at time $t$, neither is clear-certified.
The protocol relies on the following definition to characterize each node’s knowledge of what other nodes may have certified.
Definition 4.
For $\gamma\geq 0$, a block ${\sf B}$ becomes $\boldsymbol{\gamma}$-potentially certified by node $i$ at time $t$, written ${\sf B}\in\mathcal{PC}_{i}^{t}(\gamma)$, if at time $t$ there is a chain $({\sf B}_{0},\dots,{\sf B},\dots,\bar{{\sf B}})$ in ${{\sf View}}_{t}^{i}$ such that
  1. 1.
    block ${\sf B}$ is at least $k$-deep in the chain $\sf Chain(\bar{{\sf B}})$, that is $h(\bar{{\sf B}})-h({\sf B})+1\geq k$;
  2. 2.
    the chain $\sf Chain(\bar{{\sf B}})$ would have been a longest chain in $i$’s view at time $t-\gamma$, meaning $h(\bar{{\sf B}})\geq h(\tilde{{\sf B}})$ for any $\tilde{{\sf B}}\in{{\sf View}}_{i}^{t-\gamma}$.
We can now state the definition of the Stubborn Nakamoto. Recall that each node $i$ maintains ${\sf Log}_{i}^{t}$, a local copy of the ledger that consists of all finalized blocks. The local copy of the ledger ${\sf Log}_{i}^{t}$ is required to agree with the local copy of the ledger ${\sf Log}_{j}^{\tau}$ of any other honest node $j$.